1/2 log(2,x^2) = log(4,4x)
Answer by stanbon(48535)

(1/2)log(2,x^2) = log(4,4x)
-------------------------------
log2 x^2^(1/2) = log4 4x
log2 x = log4 4x
log2 x = 2log2 4x
log2 x = log2 (4x)^2
log2 x = log2 (16x^2)
x = 16x^2
16x^2 -x = 0
x(16x - 1) = 0
x = 0 or x = 1/16
---------------------------
Checking those "answers":
Does x=0 ?
log2 x = log2 (16x^2)
log2 0 = log2 0
OK
-------------
Does x = 1/16 ?
log2 (1/16) = log2 (16(1/16)^2)
log2 (1/16) = log2 (1/16)
OK
====================================
Cheers,
Stan H.